OpenStudy (zmudz):

The function \(f\) satisfies \(f(x) + f(2x + y) + 5xy = f(3x - y) + 2x^2 + 1\) for all real numbers \(x\), \(y\). Determine the value of \(f(10)\).

OpenStudy (misty1212):

can't seem to hit on the right combination

OpenStudy (zarkon):

x=10,y=5

OpenStudy (zarkon):

in general x=2y or y=x/2 then you can actually find f(x)

OpenStudy (jhannybean):

How did you find x = 2y?

OpenStudy (zarkon):

2x+y=3x-y then 2y=x